# Strong-to-weak spontaneous breaking of 1-form symmetry and intrinsically mixed topological order

### v2

@@ Appendix F Proof of exact ’t Hooft symmetry
Since each $h_{e}$ are independently drawn from a bimodal distribution $h_{e}=\pm h$ or Gaussian distribution with zero mean, $-h_{e}$ follows the same probability distribution as $h_{e}$. Therefore, the Hamiltonian $H(\{h_{e}^{X}\})$ describes the same disordered system, and the two density matrices in Eq. (156) are hence equal to each other.
@@ G.1 ST-SSB phase
[…]
where $H(s)=H_{0}+s\sum_{e}h_{e}Z_{e}$ so $\partial_{s}H(s)=\sum_{e}h_{e}Z_{e}$. We assume the presence of a gap $\Delta E$ for all $s\in[0,1]$ for the TO phase. $F(\Delta Et)$ is an antisymmetric function, so that $D(s)$ is Hermitian. Furthermore, it must be maximally quickly decaying given that its Fourier transform $\hat{F}(\omega)=-\frac{1}{\omega}$ for $|\omega|\geq 1$. It can be shown that $F(\Delta Et)$ can be chosen to be (almost) exponentially decaying for $\Delta Et>1$. This guarantees that $D(s)$ is local.  [^21] Specifically, it is local because we can split the integral over $-\frac{1}{\Delta E}\leq t\leq\frac{1}{\Delta E}$ and $\frac{1}{\Delta E}\leq|t|\leq\infty$. Then the first integral is bounded due to Lieb-Robinson bounds and the second integral is bounded due to decay of $F(\Delta Et)$.
We can commute the $B_{p}$ operators through $\partial_{s}H(s)$ so we can replace $H(s)$ by $-\lambda_{A}\sum A_{v}+s\sum_{e}h_{e}Z_{e}$. Clearly, $D(s)$ commutes with $\{B_{p}\}$, so $U$ commutes with $\{B_{p}\}$. The density matrix is given by
@@
[…]
@@
where the sum is over links that have an endpoint on the vertex $v$. The above can be obtained by trotterizing the time evolution to alternate between evolving with $s\sum h_{e}Z_{e}$ and $-\lambda_{A}\sum A_{v}$, and then rearranging terms appropriately. Note that flipping $h_{e}\to-h_{e}$ does not only take $\partial_{s}H_{s}\to-\partial_{s}H_{s}$ but also modifies $\tilde{H}_{A}(s,t^{\prime})$. Therefore, it does not simply take $D(s)\to-D(s)$ and $U\to U^{\dagger}$.
@@
[…]
@@
[…]
where we dropped terms in the expansion of the exponential that are even in $\tilde{t}$ because $F(\tilde{t})$ is an odd function of $\tilde{t}$. Using $\Delta E=4\lambda_{A}$ and writing the sine in terms of exponentials, we get
@@ Appendix H Scaling of $O_{1,z}$ for Gaussian disorder
[…]
@@
If $\frac{\lambda_{B}}{\Delta_{h}}\ll 1$, then the leading contribution comes from the strings of length 1, i.e. the plaquettes at the boundary of $\gamma$. It is very unlikely for $B_{p}$ to overpower $\Delta_{h}$ on two consecutive plaquettes, so in any particular disorder realization the addition of $-\lambda_{B}\sum B_{p}$ to the Hamiltonian only leads to modification of the ground state by some pairs of nearest neighbor flips of $B_{p}$ (from a single $h_{e}$ getting overpowered). If these nearest neighbors happen to occur at the boundary of $\gamma$, then the sign of $\langle W(\gamma)\rangle$ would change. From the above reasoning, to leading order, we can ignore the bulk of $\gamma$, and only consider plaquettes at the boundary of $\gamma$. Originally, the probability to have $B_{p}=\pm 1$ are given by
@@
with the addition of the $-\lambda_{B}\sum B_{p}$, we have
@@
where $\epsilon=\frac{\sqrt{2}\lambda_{B}}{\sqrt{\pi}\Delta_{h}}$. The change in $\langle W(\gamma)\rangle$ is given by the imbalance of $\prod_{p\in\partial\gamma}B_{p}$ to be $+1$ rather than $-1$. When $\lambda_{B}=0$, the probability for $+1$ is the same as the probability for $-1$. When $\lambda_{B}\neq 0$, this probability is
@@
Note that the above makes it seem like this fails for $\sqrt{\pi}\Delta_{h}<2\sqrt{2}\lambda_{B}$, but this is only because we have expanded the error function in the limit $\lambda_{B}\ll\Delta_{h}$. If we used the exact error function rather than this approximation, then the quantity inside the parenthesis above is always $<1$, so we always get exponential decay in $|\partial\gamma|$.

### v3

@@ Appendix F Proof of exact ’t Hooft symmetry
Since each $h_{e}$ are independently drawn from a bimodal distribution $h_{e}=\pm h$ or Gaussian distribution with zero mean, $h_{e}$ follows the same probability distribution as $h_{e}$. Therefore, the Hamiltonian $H(\{h_{e}^{X}\})$ describes the same disordered system, and the two density matrices in Eq. (156) are hence equal to each other.
@@ G.1 ST-SSB phase
where $U=\mathcal{P}e^{i\int_{0}^{1}dsD(s)}$ and
$D(s)=-i\int dtF(\Delta Et)e^{iH(s)t}\partial_{s}H(s)e^{-iH(s)t}$ (160)
where $H(s)=H_{0}+s\sum_{e}h_{e}Z_{e}$ so $\partial_{s}H(s)=\sum_{e}h_{e}Z_{e}$. We assume the presence of a gap $\Delta E$ for all $s\in[0,1]$ for the TO phase. $F(\Delta Et)$ is an antisymmetric function, so that $D(s)$ is Hermitian. Furthermore, it must be maximally quickly decaying given that its Fourier transform $\hat{F}(\omega)=-\frac{1}{\omega}$ for $|\omega|\geq 1$. It can be shown that $F(\Delta Et)$ can be chosen to be (almost) exponentially decaying for $\Delta Et>1$. This guarantees that $D(s)$ is local.  [^21] Specifically, it is local because we can split the integral over $\frac{1}{\Delta E}\leq t\leq\frac{1}{\Delta E}$ and $\frac{1}{\Delta E}\leq|t|\leq\infty$. Then the first integral is bounded due to Lieb-Robinson bounds and the second integral is bounded due to decay of $F(\Delta Et)$.
We can commute the $B_{p}$ operators through $\partial_{s}H(s)$ so we can replace $H(s)$ by $\lambda_{A}\sum A_{v}+s\sum_{e}h_{e}Z_{e}$. Clearly, $D(s)$ commutes with $\{B_{p}\}$, so $U$ commutes with $\{B_{p}\}$. The density matrix is given by
@@
$\begin{split}U_{Z}(s,t)&=e^{is\sum_{e}h_{e}Z_{e}t}\\ U_{A}(s,t)&=\mathcal{T}e^{-i\int_{0}^{t}dt^{\prime}\tilde{H}_{A}(t^{\prime})}\end{split}$ (163)
@@
where the sum is over links that have an endpoint on the vertex $v$. The above can be obtained by trotterizing the time evolution to alternate between evolving with $s\sum h_{e}Z_{e}$ and $\lambda_{A}\sum A_{v}$, and then rearranging terms appropriately. Note that flipping $h_{e}\to-h_{e}$ does not only take $\partial_{s}H_{s}\to-\partial_{s}H_{s}$ but also modifies $\tilde{H}_{A}(s,t^{\prime})$. Therefore, it does not simply take $D(s)\to-D(s)$ and $U\to U^{\dagger}$.
@@
$\begin{split}D(s)&\sim-i\int_{-\infty}^{\infty}dtF(\Delta Et)e^{-i\lambda_{A}\sum A_{v}t}\left(\sum h_{e}Z_{e}\right)e^{i\lambda_{A}\sum A_{v}t}\\ &=-i\int_{-\infty}^{\infty}\frac{d\tilde{t}}{\Delta E}F(\tilde{t})\sum h_{e}Z_{e}e^{\frac{2\lambda_{A}i}{\Delta E}(A_{v}+A_{v^{\prime}})\tilde{t}}\end{split}$ (165)
@@
$\begin{split}D(s)&=-\frac{i}{\Delta E}\sum h_{e}Z_{e}\int_{-\infty}^{\infty}d\tilde{t}F(\tilde{t})i\cos\left(\frac{2\lambda_{A}}{\Delta E}\right)\sin\left(\frac{2\lambda_{A}}{\Delta E}\right)(A_{v}+A_{v^{\prime}})\\ &=-\frac{i}{\Delta E}\sum h_{e}Z_{e}(A_{v}+A_{v^{\prime}})\int_{-\infty}^{\infty}d\tilde{t}F(\tilde{t})\frac{i}{2}\sin\left(\frac{4\lambda_{A}t}{\Delta E}\right)\end{split}$ (166)
where we dropped terms in the expansion of the exponential that are even in $\tilde{t}$ because F(t~)F($\tilde${t}) is an odd function of $\tilde{t}$. Using $\Delta E=4\lambda_{A}$ and writing the sine in terms of exponentials, we get
@@ Appendix H Scaling of $O_{1,z}$ for Gaussian disorder
for all $B_{p}$. This comes from the fact that the distribution of $\{h_{e}\}$ have zero mean. Now we turn a small $B_{p}$ term, so that $\frac{\lambda_{B}}{\Delta_{h}}\ll 1$. This gives a small probability for the $B_{p}$ term to overpower the $Z_{e}$ terms for a given plaquette. Specifically, for a particular disorder realization, if the $Z_{e}$ terms in a plaquette favor $B_{p}=-1$ but there is a link with $h_{e}<\lambda_{B}$, then it would be energetically favorable for $B_{p}=-1$ to flip to $B_{p}=+1$ after the addition of the $\lambda\sum B_{p}$ term to the Hamiltonian. The probability for a $B_{p}$ term to overpower the $Z_{e}$ terms is $\sim\frac{\sqrt{2}\lambda_{B}}{\sqrt{\pi}\Delta_{h}}$ as can be obtained from integrating over the Gaussian distribution $\frac{1}{\sqrt{2\pi}\Delta_{h}}\int_{-\lambda_{B}}^{\lambda_{B}}dh\ e^{-h^{2}/2\Delta_{h}^{2}}$ and expanding at small $\lambda_{B}/\Delta_{h}$.
@@
If $\frac{\lambda_{B}}{\Delta_{h}}\ll 1$, then the leading contribution comes from the strings of length 1, i.e. the plaquettes at the boundary of $\gamma$. It is very unlikely for $B_{p}$ to overpower $\Delta_{h}$ on two consecutive plaquettes, so in any particular disorder realization the addition of $\lambda_{B}\sum B_{p}$ to the Hamiltonian only leads to modification of the ground state by some pairs of nearest neighbor flips of $B_{p}$ (from a single $h_{e}$ getting overpowered). If these nearest neighbors happen to occur at the boundary of $\gamma$, then the sign of $\langle W(\gamma)\rangle$ would change. From the above reasoning, to leading order, we can ignore the bulk of $\gamma$, and only consider plaquettes at the boundary of $\gamma$. Originally, the probability to have $B_{p}=\pm 1$ are given by
@@
with the addition of the $\lambda_{B}\sum B_{p}$, we have
@@
where $\epsilon=\frac{\sqrt{2}\lambda_{B}}{\sqrt{\pi}\Delta_{h}}$. The change in $\langle W(\gamma)\rangle$ is given by the imbalance of $\prod_{p\in\partial\gamma}B_{p}$ to be +1+1 rather than 1-1. When $\lambda_{B}=0$, the probability for +1+1 is the same as the probability for 1-1. When $\lambda_{B}\neq 0$, this probability is
@@
Note that the above makes it seem like this fails for $\sqrt{\pi}\Delta_{h}<2\sqrt{2}\lambda_{B}$, but this is only because we have expanded the error function in the limit $\lambda_{B}\ll\Delta_{h}$. If we used the exact error function rather than this approximation, then the quantity inside the parenthesis above is always <1<1, so we always get exponential decay in $|\partial\gamma|$.